OpenStudy (anonymous):

a graph of a parabola has a line of symmetry x=3 and contains points (1,0) and (4,-3). determinean equation for the parabola.

OpenStudy (anonymous):

let the eq of parabola be (x-3)^2=A(y-B) now we need to determine A and B

OpenStudy (anonymous):

by puttin 1,0 we have AB =-4 and 1=-A(3+B) solving the equation B=-AB(3+B) or B=+12 +4B or B=-4 and A=1 hence the reqd parabola is (x-3)^2=(y+4)